# Enhanced premelting at the ice–rubber interface using all-atom molecular dynamics simulation

### v1

@@ 1 Introduction
The ice–rubber interface plays a key role in many industrial applications, including vehicle tires and shoe outsoles. For tires, rubber materials help reduce the risk of slipping and allow efficient transmission of engine power and braking force to icy road surfaces. Rubber friction arises from its viscoelasticity and surface adhesion on solid surfaces, typically showing a bell-shaped dependence on sliding velocity 1, 2, 3. In contrast, the tribology of rubber on ice differs markedly because of the unique properties of ice. Ice is exceptionally slippery, and its low friction has been attributed to several mechanisms, including frictional heating, premelting layers, pressure melting, and the formation of thin liquid-like surface films 4, 5, 6, 7, 8, 9. Nevertheless, the fundamental origin of this slipperiness remains unresolved.
@@
A thin liquid-like layer, known as the premelting layer, forms on ice surfaces and strongly affects interfacial behavior 10, 11, 12. Its thickness has been reported from a few nanometers to micrometers, and although debated, it is widely accepted that the layer behaves like a liquid. Near the melting point, it is highly sensitive to temperature, where changes of only a few kelvins can strongly alter its morphology, structure, and dynamics 13, 14, 15, 16. Molecular simulations have been widely used to probe the structure and dynamics of premelting layers 17, 18, 14, 19, 20, 21, 16. Recent computational studies show that small molecules can locally alter ice surfaces. For example, ions disrupt the local ice structure 22, 23, and short hydrophilic polymers induce partial melting and suppress ice growth 24. However, the effects of polymer interfaces remain largely unexplored.
@@
The ice–polymer (or rubber) interface has been studied experimentally for decades, demonstrating that the surface properties of ice are strongly influenced by the chemistry and microstructure of the contacting material. Friction measurements showed that the viscosity of the premelting film is sensitive to surface chemistry 25. Hydrophobic coatings reduce ice adhesion, whereas hydrophilic polymers interact with liquid-like films to form hydration layers 26, 27. Lecadre et al.28 suggested that friction is governed by the viscoelasticity of contacting materials at low temperatures, but dominated by premelting layers at higher temperatures. Molecular dynamics simulations further revealed that wettability strongly affects premelting-layer friction29, with similar behavior seen in supercooled water 30. More realistic interface models, including polymer and graphene interfaces, showed that hydrogen bonding with the contacting surface alters the orientation, structure, and dynamics of water molecules in the premelting layers 31. These findings suggest that both the chemical composition and nanostructure of rubber chains can modify premelting layers, producing behaviors distinct from those at the pure ice–vapor interface. However, the atomistic details of such rubber-induced effects remain largely unexplored.
@@
Rubber is mainly composed of nonpolar polymers such as styrene and butadiene, often combined with cross-linkers and fillers. Its hydrophobic nature is expected to affect the ice surface through van der Waals interactions. In this study, we use molecular dynamics simulations to investigate the atomistic structure and dynamics at the ice–rubber interface, as slipperiness is closely linked to molecular mobility 32. We find that rubber enhances structural disorder in the premelting layers while suppressing their dynamics due to confinement. Importantly, we also show that the dynamics of interfacial rubber chains change cooperatively with premelting water molecules, and at near the melting temperature, a mixed rubber–water layer emerges.
2 Methods
@@ 2.1 Molecular Dynamics Simulation
Figure 1: Simulation system. (A) Chemical structure of styrene-butadiene rubber used in this work. The number of monomer unit is indicated and sequence of monomer was randomly shuffled to . (B) Representative snapshot of the system. The top side is the ice–rubber interface, while the lower is the ice–vapor interface.
@@ 2.3 Diffusion coefficient
The molecular dynamics of premelting layers was evaluated using the mean square displacement (MSD),
@@
where $\langle\cdot\rangle$ means the ensemble average, and $\bm{r}(t)$ is the position of the oxygen atom. In this calculation, we counted molecules that stayed in the combined region of layers 1 and 2 during the measurement time, $\Delta$. The diffusion coefficient, $D$, was computed from the sloop of MSD in the nearly linear region (after lag time 100 ps),
$D=\frac{1}{6}\frac{d}{dt}\mathrm{MSD}(t)$ (2)
2.4 Analysis of single molecular dynamics using machine learning
Although MSD is useful for characterizing system-averaged molecular dynamics, single-molecule dynamics cannot be fully captured. To distinguish liquid-like mobile molecules from solid-like immobile molecules, a machine-learning approach for single-molecular dynamics analysis has been developed 44, 45, 46. In this work, we employ this method to clarify the rubber-inducing effect on the premelting molecules at single-molecule resolution. In the machine learning approach, we first sample the short-term trajectories from the rubber-ice interface, $\bm{x}$, and those from bulk ice systems, $\bm{x^{\prime}}$. Any type of trajectories can be used for $\bm{x^{\prime}}$, but we employed a time series of the oxygen atom displacement in the water molecules for 64 ps, according to our previous work 46. Both $\bm{x}$ and $\bm{x^{\prime}}$ are input into a neural network, which outputs a score, $g(\bm{x})$, that quantifies the difference from the solid bulk water system,
@@ 2.4 Analysis of single molecular dynamics using machine learning
where $\mathbb{E}$ represents the expectation value taken over samples $\bm{x}$ sampled from the distribution of $\bm{y}$, and $f^{*}(\bm{x}^{\prime})$ is the function represented by the neural network, which is a simple multi-layer perceptron. The neural network model trained in ref. 46 was used. Higher values of $g(\bm{x})$ correspond to molecular dynamics that differ from the solid phase. Following our previous work, a threshold value of $g(\bm{x})$, which was determined based on the comparison of liquid and solid bulk systems, was used to distinguish liquid and solid molecules in the premelting layers.
2.5 Analysis of polymer dynamics
@@ 3.1 Enhanced disorder of premelting layers at the ice–rubber interface
As premelting layers are highly sensitive to temperature near the melting point 13, 14, we compare the atomistic surface structures at 264 and 269 K for the ice–vapor and ice–rubber interfaces. Snapshots show that the surface structures are similar at 264 K, whereas at 269 K, the third layer becomes more disordered at the rubber interface, compared to the vapor interface (Fig. 2A). The density profiles show that rubber and premelting layers coexist in a very thin surface layer at both temperatures (Fig. 2B). However, the density profiles at 269 K show weaker peaks at the rubber interface compared to the vapor interface, while their profiles at 264 K are nearly identical (Fig. 2C). This difference is evident even in the 3rd ice layer, which is not directly exposed to rubber atoms. Interestingly, at 269 K, the third layer at the ice–rubber interface exhibits comparable peak heights and widths in the density profile to the second layer at the ice–vapor interface, indicating structural similarity between the two layers (Fig. 2C). Comparison of the profiles at different temperatures in the range of 254–269 K (Fig. S1) indicates a sharp increase in ice premelting between 267 and 269 K, accompanied by enhanced penetration of rubber atoms into the premelting layer.
@@
Figure 3: Comparison of the dynamical properties of ice–rubber and ice–vapor interfaces. (A) Mean squared displacement (MSD) of premelting layers (combined Layer 1 and Layer 2) in the ice–rubber interface. $\alpha$ is the scaling exponent of MSD. (B) Diffusion coefficient as a function of temperature. (C) Distribution of the score, $g(\bm{x})$, which quantifies deviation from bulk solid behavior in single-molecule short-term dynamics. Layer 1–4 represent the distributions from the ice–rubber interface at 269 K, while ”Ice” and ”Water” correspond to bulk solid and liquid systems, respectively. The black dashed line indicates the threshold used to classify molecules as solid-like or liquid-like. (D–F) Ratio of liquid-like molecules in Layers 1–3 at different temperatures. Error bar shows standard deviation of the mean values computed using 5 ns time-blocks in (B) and (C–F).
@@ 3.2 Suppressed water dynamics by rubber confinement
Building on the observation of enhanced structural disorder at the ice–rubber interface, we use dynamical properties to clarify whether the premelting layers behave as mobile liquid. Given that the relaxation times of the first and second layers are on the order of several tens of picoseconds 18, 46, we treated these two layers as a single region and computed the MSD of molecules within the combined region. For the rubber interface, MSD increased significantly between 264 and 269 K (Fig. 3A). The scaling exponent, $\alpha$ ($\rm{MSD}\propto t^{\alpha}$), ranged in 0.85–0.89 for the vapor interface, and 0.77–0.87 for the rubber interface (Tab. S1), indicating the glassy behavior of the premelting layers 19. The ice–rubber interface showed a lower $\alpha$ value compared to the ice–vapor interface, suggesting that the presence of rubber molecules enhances the confinement of water molecules within the premelting layers.
@@
To further clarify the temperature dependence of interfacial dynamics, we computed the diffusion coefficients for the ice–vapor and ice–rubber interfaces. As shown in Fig. 3B, the diffusion coefficient at the ice–vapor interface increases linearly with temperature, whereas the ice–rubber interface exhibits a change in the slope between 264 and 267 K. These results highlights distinct differences in both diffusivity and temperature dependence between the two interfaces.
@@
Besides the mean behavior within the layer, we speculate that the local nanostructure of the rubber molecules contributes to the heterogeneity in the layer, thus requiring a single-molecular dynamics analysis. Here, from the measurement of the short-term molecular dynamics (64 ps) of ice premelting layers at the single-molecule resolution 46, we quantify the ratio of mobile-liquid-like molecules. We computed the score $g(\bm{x})$ for the four premelting layers, which represents differences from the bulk ice molecules on the short-term dynamics, and compared the score to those of bulk solid and liquid water systems. For the rubber interface, Layers 1 and 2 show similar properties to the bulk liquid system (Fig. 3C). The threshold to classify solid and liquid molecules was determined at 269 K (Fig. 3C) and subsequently applied to all other temperatures, following the approach used in previous work 46.
@@
Figure 4: Analysis of the atomistic structure of the ice–rubber interface. (A) Snapshot of Layer 1 at 269 K, showing water molecules as van der Waals spheres and rubber atoms as a surface representation. (B) Root-mean-square fluctuation (RMSF) of surface rubber atoms in Layer 1. Error bar shows standard deviation of the mean values computed using 5 ns time-blocks. (C) Representative snapshot of a region within Layer 1, where the densities of water and rubber are comparable, at 254 and 267 K. Water molecules are shown as oxygen atom positions and are colored by the score $g(\bm{x})$, which quantifies deviation from bulk ice. Higher values (yellow) indicate greater mobility. Rubber atoms are shown as green spheres.

### v2

@@ 1 Introduction
The ice–rubber interface plays a key role in many industrial applications, including vehicle tires and shoe outsoles. For tires, rubber materials help reduce the risk of slipping and allow efficient transmission of engine power and braking force to icy road surfaces. Rubber friction arises from its viscoelasticity and surface adhesion on solid surfaces, typically showing a bell-shaped dependence on sliding velocity grosch1963relation, ROBERTS197675, chernyak1986theory. In contrast, the tribology of rubber on ice differs markedly because of the unique properties of ice. Ice is exceptionally slippery, and its low friction has been attributed to several mechanisms, including frictional heating, premelting layers, pressure melting, and the formation of thin liquid-like surface films southern1972friction, persson2015ice, rosenberg2005ice, kietzig2010physics, tuononen2016multiscale, hemette2021thermal. In recent years, systematic experiments have revealed that the interfacial temperature affects the thickness of the viscoelastic ice surface film canale2019nanorheology, that the friction mode changes depending on surface temperature and pressure liefferink2021friction, and that surface chemistry modulates molecular diffusion and, consequently, friction canale2019nanorheology, baran2022ice. Despite these advances in understanding ice tribology, the microscopic origins of ice–rubber interfacial tribology remain poorly understood.
@@
A thin liquid-like layer, known as the premelting layer, forms on ice surfaces and strongly affects interfacial behavior dash2006physics, li2007surface, slater2019surface. Its thickness has been reported from a few nanometers to micrometers, and although debated, it is widely accepted that the layer behaves like a liquid. Near the melting point, it is highly sensitive to temperature, where changes of only a few kelvins can strongly alter its morphology, structure, and dynamics sazaki2012quasi, sanchez2017experimental, nagata2019surface, llombart2020surface. Molecular simulations have been widely used to probe the structure and dynamics of premelting layers, supporting their liquid-like nature conde2008thickness, pfalzgraff2011comparative, sanchez2017experimental, kling2018structure, pickering2018grand, qiu2018so, llombart2020surface. Recent computational studies have shown that small molecules can locally alter ice surfaces. For example, ions disrupt the local ice structure hudait2017sink, berrens2022effect, and hydrophilic polymers of low molecular weight induce partial melting and suppress ice growth bachtiger2021atomistic. These computational studies have provided molecular insights into ice–vapor interfaces.
@@
The ice–polymer (or rubber) interface has been studied experimentally for decades, demonstrating that the surface properties of ice are strongly influenced by the chemistry and microstructure of the contacting material. Friction measurements showed that the viscosity of the premelting film is sensitive to surface chemistry canale2019nanorheology. Hydrophobic coatings reduce ice adhesion, whereas hydrophilic polymers interact with liquid-like films to form hydration layers chen2017icephobic, pallbo2024enhanced. Lecadre et al.lecadre2020ice suggested that friction is governed by the viscoelasticity of contacting materials at low temperatures, but dominated by premelting layers at higher temperatures. Molecular dynamics simulations further revealed that wettability strongly affects premelting-layer frictionbaran2022ice, with similar behavior seen in supercooled water gasparotto2022interfaces. More realistic interface models, including polymer and graphene interfaces, showed that hydrogen bonding with the contacting surface alters the orientation, structure, and dynamics of water molecules in the premelting layers skountzos2024interfacial. These findings suggest that both the chemical composition and nanostructure of rubber chains can modify premelting layers, producing behaviors distinct from those at the pure ice–vapor interface. However, the atomistic details of such rubber-induced effects remain largely unexplored.
@@
Rubber is mainly composed of nonpolar polymers such as styrene and butadiene, often combined with cross-linkers and fillers. Its hydrophobic nature is expected to affect the ice surface through Van der Waals interactions. In this study, we use molecular dynamics simulations to investigate the atomistic structure and dynamics at the ice–rubber interface, as slipperiness is closely linked to molecular mobility weber2018molecular. We find that rubber enhances structural disorder in the premelting layers while suppressing their dynamics due to confinement. Importantly, we also show that the dynamics of interfacial rubber chains change cooperatively with premelting water molecules, and at near the melting temperature, a mixed rubber–water layer emerges.
2 Materials and Methods
@@ 2.1 Molecular Dynamics Simulation
Figure 1: Simulation system. (A) Chemical structure of styrene-butadiene rubber used in this work. We generated a single chain by randomly shuffling the monomers, using the indicated number of monomer units. (B) Representative zoomed-in snapshot of the interface, with the full simulation system shown in Fig. S1. The top side is the ice–rubber interface, while the lower is the ice–vapor interface.
@@ 2.3 Diffusion coefficient
Molecular dynamics of premelting layers was evaluated using the mean square displacement (MSD) along the directions parallel to the interface,
@@
where $\langle\cdot\rangle$ means the ensemble average, and $\bm{r}(t)$ is the position of the oxygen atom. In this calculation, we counted molecules that remained within the specified region for the measurement time, $\Delta$: (1) Layer 1 and (2) the combined region of Layers 1 and 2. A comparison of these definitions is shown in the Results section. The parallel diffusion coefficient kling2018structure, louden2018ice, baran2022ice, $D$, was computed from the slope of MSD in the nearly linear region (after lag time 100 ps),
$D=\frac{1}{2d}\frac{d}{dt}\mathrm{MSD}(t)$ (2)
where $d$ is the number of dimension and, for the parallel diffusion $d=2$.
2.4 Analysis of single molecule dynamics using machine learning
Although MSD is useful for characterizing system-averaged molecular dynamics, single-molecule dynamics cannot be fully captured. Although bond-order parameters have been widely used to detect local ice structures pickering2018grand, mochizuki2023microscopic, these descriptors rely on instantaneous local structures. However, the polymer–ice interfaces are highly heterogeneous, making it difficult to judge the local states based on single-frame structural snapshots. Therefore, to distinguish liquid-like mobile molecules from solid-like immobile molecules, a machine-learning approach for single-molecular dynamics analysis has been developed endo2019detection, yasuda2023combining, yasuda2024layer. In this work, we employed this method to clarify the rubber-induced effect on the premelting molecules at single-molecule resolution. In the machine learning approach, we first sample the short-term trajectories from the rubber-ice interface, $\bm{x}$, and those from bulk ice systems, $\bm{x^{\prime}}$. Any type of trajectories can be used for $\bm{x^{\prime}}$, but we employed a time series of the oxygen atom displacement in the water molecules for 64 ps, according to our previous work yasuda2024layer. Both $\bm{x}$ and $\bm{x^{\prime}}$ are input into a neural network, which outputs a score, $g(\bm{x})$, that quantifies the difference from the solid bulk water system,
@@ 2.4 Analysis of single molecule dynamics using machine learning
where $\mathbb{E}$ represents the expectation value taken over samples $\bm{x}$ sampled from the distribution of $\bm{y}$, and $f^{*}(\bm{x}^{\prime})$ is the function represented by the neural network, which is a simple multi-layer perceptron. The neural network model trained in ref. yasuda2024layer was used. Higher values of $g(\bm{x})$ correspond to molecular dynamics that differ from the solid phase. Following our previous work, a threshold value of $g(\bm{x})$, which was determined based on the comparison of liquid and solid bulk systems, was used to distinguish liquid and solid molecules in the premelting layers.
2.5 Analysis of polymer fluctuations
@@ 3.1 Enhanced disorder of premelting layers at the ice–rubber interface
As premelting layers are highly sensitive to temperature near the melting point sazaki2012quasi, sanchez2017experimental, we compare the atomistic surface structures at 264 and 269 K for the ice–vapor and ice–rubber interfaces. Snapshots show that the surface structures are similar at 264 K, whereas at 269 K, the third layer becomes more disordered at the rubber interface, compared to the vapor interface (Fig. 2A). The density profiles show that rubber and premelting layers coexist in a very thin surface layer at both temperatures (Fig. 2B). However, the density profiles at 269 K show weaker peaks at the rubber interface compared to the vapor interface, while their profiles at 264 K are nearly identical (Fig. 2C). This difference is evident even in the 3rd ice layer, which is not directly exposed to rubber atoms. Interestingly, at 269 K, the third layer at the ice–rubber interface exhibits comparable peak heights and widths in the density profile to the second layer at the ice–vapor interface, indicating structural similarity between the two layers (Fig. 2C).
@@
Comparison of the profiles at different temperatures in the range of 254–269 K (Fig. S2) indicates a sharp increase in ice premelting between 264 and 267 K, accompanied by enhanced penetration of rubber atoms into the premelting layer. To further support the disruption of the hydrogen-bond network, we computed the orientational order parameter (Fig. S3). The reduction in peak intensity indicates rubber-induced orientational disordering above 264 K in Layers 2–4 (Fig. S4), in consistent with the density profiles.
@@ 3.2 Suppressed water dynamics by rubber confinement
Building on the observation of enhanced structural disorder at the ice–rubber interface, we use dynamical properties to clarify whether the premelting layers behave as mobile liquid. MSD was computed for water molecules that remained in Layer 1 at the ice–vapor and at ice–rubber interfaces. For the rubber interface, MSD increased significantly between 264 and 269 K (Fig. 3A). The scaling exponent of MSD, $\alpha$ ($\rm{MSD}\propto t^{\alpha}$), is 1 for Brownian motion, and if the exponent is lower than 1, the dynamics are subdiffusive, which is typically observed in glassy or viscoelastic materials and crowding environments kling2018structure. The scaling exponent ranged in 0.92–0.96 for the vapor interface, and 0.77–0.85 for the rubber interface (Tab. S1), indicating the glassy behavior of the premelting layers kling2018structure. The ice–rubber interface showed a lower $\alpha$ value compared to the ice–vapor interface, suggesting that the presence of rubber molecules enhanced the confinement of water molecules within the premelting layers.
Figure 3: Comparison of the dynamical properties of ice–rubber and ice–vapor interfaces. (A) Mean squared displacement (MSD) of premelting layers (combined Layer 1 and Layer 2) in the ice–rubber interface. $\alpha$ is the scaling exponent of MSD. (B) Parallel diffusion coefficient as a function of temperature. Bulk liquid values at 0 MPa were taken from Baran et al. baran2023self. Values of the parallel diffusion coefficients are listed in Tab. S1. (C) Distribution of the score, $g(\bm{x})$, which quantifies deviation from bulk solid behavior in single-molecule short-term dynamics. Layer 1–4 represent the distributions from the ice–rubber interface at 269 K, while ”Ice” and ”Water” correspond to bulk solid and liquid systems, respectively. The black dashed line indicates the threshold used to classify molecules as solid-like or liquid-like. (D–F) Ratio of liquid-like molecules in Layers 1–3 at different temperatures. Error bars represent the standard error of the mean calculated using time block averaging (5 ns per block) in (B) and (C–F).
@@
To further clarify the temperature dependence of interfacial dynamics, we computed the parallel diffusion coefficients for the ice–vapor and ice–rubber interfaces. As shown in Fig. 3B, the parallel diffusion coefficient at the ice–vapor interface increased linearly with temperature, whereas the ice–rubber interface exhibited a change in the slope between 264 and 267 K. These results highlight distinct differences in both diffusivity and temperature dependence between the two interfaces. Furthermore, our results demonstrate higher the parallel diffusion coefficients at the ice–vapor interface than the diffusion coefficient of bulk liquid, consistent with previous studies kling2018structure, louden2018ice, whereas the rubber interface exhibited suppressed dynamics (Fig. 3B). In contrast, a previous study reported enhanced dynamics at smooth hydrophobic interfaces baran2022ice. The suppressed dynamics in our atomistic rubber interfaces showed suggests that the structural roughness of the interface induced confinement effects.
We compared the parallel diffusion coefficients at the rubber interface for (1) Layer 1 alone and (2) the combined Layers 1 and 2. At most temperatures, Layer 1 alone showed higher diffusion coefficients than the combined Layers 1 and 2. However, at 269 K, the combined Layers 1 and 2 showed a higher diffusion coefficient. We interpret this as the result of polymer-induced confinement within Layer 1, while Layer 2 provides a partially melted ice structure that facilitates water diffusion (Fig. S5).
While these simulations were performed at $1~\mathrm{MPa}$, the density profiles at 254 and 269 K are nearly the same at $0.1~\mathrm{MPa}$, indicating that the pressure effect is negligible within this range (Fig. S6). Similarly, the difference in the diffusion coefficient between 0.1 MPa and 1 MPa was minor (Fig. S7).
3.3 Liquid-like molecular ratio for each layer
@@
Besides the mean behavior within the layer captured by MSD, we speculate that the local nanostructure of the rubber molecules contributes to the heterogeneity in the layer, thus requiring a single-molecular dynamics analysis. Here, from the measurement of the short-term molecular dynamics (64 ps) of ice premelting layers at the single-molecule resolution yasuda2024layer, we quantify the ratio of mobile-liquid-like molecules. This approach assigns to each short-term single-molecule trajectory, $\bm{x}$, a score, $g(\bm{x})$, that quantifies how different its dynamics are from those observed in the bulk ice. The model was trained on bulk solid and bulk liquid water at 269 K. Despite being trained only on bulk systems at a single temperature, the model successfully generalizes to interface systems, including different ice crystal faces across different temperatures yasuda2024layer. In our ice–rubber interfaces, this is evidenced by the gradual shifts in $g(\bm{x})$ distributions with temperature (Fig. S8), demonstrating the robustness of the approach. Focusing on 269 K, we computed the score $g(\bm{x})$ for the four premelting layers, and compared the score to those of bulk solid and liquid water systems. For the rubber interface, Layers 1 and 2 show similar properties to the bulk liquid system (Fig. 3C), indicating that these layers predominantly consist of liquid-like molecules.
The threshold for distinguishing liquid-like from solid-like molecules was optimized at 269 K to achieve balanced classification accuracy for both solid and liquid molecules yasuda2024layer. We applied this threshold consistently across all temperatures. To validate the threshold at the ice–rubber and ice–vapor interfaces, we examined Layer 4, which serves as a reference region consisting predominantly of solid-like molecules. Specifically, the liquid ratio in Layer 4 was about 6% at the ice-vapor interface, and 8% at the ice-rubber interface at 254 K (Fig. S9), which are sufficiently low values. At the higher temperatures, it increased to about 15% at the ice-vapor interface, which is the similar liquid ratio as obtained in our previous work yasuda2024layer. The difference in the liquid ratio between the bulk solid and Layer 4 at the ice–vapor interfaces at 269 K is attributed to both the uncertainty of the model and interface-enhanced mobility. Considering these points, the threshold reasonably captures the differences between solid–like and liquid-like dynamics, as a semi-qualitative measurement.
@@
Figure 4: Analysis of the atomistic structure of the ice–rubber interface. (A) Snapshot of Layer 1 at 269 K, showing water molecules as Van der Waals spheres and rubber atoms as a surface representation. (B) Root-mean-square fluctuation (RMSF) of surface rubber atoms in Layer 1. Error bars represent the standard error of the mean calculated using time block averaging (5 ns per block). (C) Representative snapshot of a region within Layer 1, where the densities of water and rubber are comparable, at 254 and 267 K. Water molecules are shown as oxygen atom positions and are colored by the score $g(\bm{x})$, which quantifies deviation from bulk ice. Higher values (yellow) indicate greater mobility. Rubber atoms are shown as green spheres.